OpenStudy (anonymous):

clear parenthesis and comebine 7+3(2a-5)-(6a-8b)-2b

OpenStudy (anonymous):

6b-8

OpenStudy (aroub):

7+3=10 ,10(2a-5)=20a-50-6a+8b-2b =14a+6b-50